Express sin^5x
as a linear combination of sin kx and coskx, k belongs to integers.
1
Expert's answer
2020-02-19T10:00:06-0500

We know

"sin3x=3sinx-4sin^3x"

"\Rightarrow sin^3x=\frac{3sinx-sin3x}{4}"

Also

"cos2x=1-2sin^2x"

"or \space sin^2x=\frac{1-cos2x}2"

We will use above two formula to write "sin^5x" as linear combination of sine and cosine


"sin^5x"

"=sin^2x\cdot sin^3x"

"=\left(\frac{1-cos(2x)}{2}\right)\left(\frac{3sin(x)-sin(3x)}{4}\right)"

"=\frac{1}{8}\left(1-cos(2x)\right)\left(3sin(x)-sin(3x)\right)"

"=\frac{1}{8}\left(1\left(3sin(x)-sin(3x)\right)-cos(2x)\left(3sin(x)-sin(3x)\right)\right)" [ multiplying

"=\frac{1}{8}\left(3sin(x)-sin(3x)-3cos(2x)sin(x)+sin(3x) cos(2x)\right)" ...............(1)

This could be the answer but it can also further expanded as below using following formula

"sin(A+B)+sin(A-B)=2sinAcosB"

"Sin(A+B)-Sin(A-B)=2cosAcosB"

From (1)

"=\frac{3}{8}sin(x)-\frac{1}{8}sin(3x)-\frac{3}{8}cos(2x)sin(x)" "+\frac{1}{8}sin(3x)cos(2x)"

"=\frac{3}{8}sin(x)-\frac{1}{8}sin(3x)-\frac{3}{16}*(2cos(2x)sin(x))" "+\frac{1}{16}(2sin(3x)cos(2x))"

"=\frac{3}{8}sin(x)-\frac{1}{8}sin(3x)-\frac{3}{16}(sin(2x+x)" "-sin(2x-x))+\frac{1}{16}(sin(3x+2x)+sin(3x-2x))"

"=\frac{3}{8}sin(x)-\frac{1}{8}sin(3x)-\frac{3}{16}sin(3x)" "+\frac{3}{16}sin(x)+\frac{1}{16}sin(5x)+\frac{1}{16}sin(x)"

"=(\frac{3}{8}+\frac{3}{16}+\frac{1}{16})sin(x)+" "(-\frac{1}{8}-\frac{3}{16})sin(3x)+\frac{1}{16}sin\left(5x\right)"

"=\frac{5}{8}sin(x)-\frac{5}{16}sin(3x)+\frac{1}{16}sin(5x)"